m is the midpoint of line JK,JM=x over 8 and JK= 3x over 4 minus 6. find MK
Answer by cleomenius(959)

JM + MK = JK.
JK - JM = MK
(3x/4) -6 - (x/8) = MK
(6x/8) -6 - (x/8) = MK
=========================
Multiply all by 8
=========================
6x - 48 - x = 8MK
5x - 48 = 8 MK
=============================
Divide all by 8
===============================
(5/8)x - 6 = MK
=============================
Cleomenius.